OpenStudy (anonymous):

help?!?!(medal will be given) Line EF has an equation of a line y = −2x + 7, and line GH has an equation of a line y = 1 over 2x −1. These two lines are parallel because the slopes of the lines are equal perpendicular because the slopes are opposite reciprocals the same line when graphed intersecting but not at a right angle

OpenStudy (anonymous):

@esshotwired @phi will the answer be the last one?

OpenStudy (phi):

what is the slope of the first equation ? what is the slope of the second equation ?

OpenStudy (anonymous):

the first one −2x the second 1 over 2x

OpenStudy (phi):

the slope is just the number part (leave out the x) if you "flip" the second slope and multiply by -1 what do you get ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

\[ \frac{1}{2} \] flip and multiply by -1 \[-1 \cdot \frac{2}{1} \] -2

OpenStudy (phi):

in other words the slopes are "negative reciprocals" when you see this, you have two lines that cross and form 90º angles

OpenStudy (anonymous):

ah so the lines are perpendicular.

OpenStudy (phi):

yes. same slope means parallel negative reciprocal slopes means perpendicular

OpenStudy (phi):

though your question calls negative reciprocal "opposite reciprocal"

OpenStudy (anonymous):

thx